# Modeling Nonlinear Oscillator Networks Using Physics-Informed Hybrid Reservoir Computing

### v1

@@ Abstract
Surrogate modeling of non-linear oscillator networks remains challenging due to discrepancies between simplified analytical models and real-world complexity. To bridge this gap, we investigate hybrid reservoir computing, combining reservoir computing with “expert” analytical models. Simulating the absence of an exact model, we first test the surrogate models with parameter errors in their expert model. Second, we assess their performance when their expert model lacks key non-linear coupling terms present in an extended ground-truth model. We focus on short-term forecasting across diverse dynamical regimes, evaluating the use of these surrogates for control applications. We show that hybrid reservoir computers generally outperform standard reservoir computers and exhibit greater robustness to parameter tuning. Notably, unlike standard reservoir computers, the performance of the hybrid does not degrade when crossing an observed spectral radius threshold. Furthermore, there is good performance for dynamical regimes not accessible to the expert model, demonstrating the contribution of the reservoir.
@@ Introduction
Networks of oscillators appear widely across engineering, and in both the physical and biological sciences. When the networks are non-linear their dynamical behavior can be complex, displaying synchronization, chaos, and traveling waves [1, 2, 3]. Creating analytical or data-driven models that can predict their dynamics is important in applications, for example in producing surrogate models. Downstream applications for a surrogate model of non-linear oscillator networks (NLONs) include smart electrical grid optimization [4, 5, 6], biological computing [7], synthetic biology [8, 9] and the diagnosis and treatment of neurological disorders such as epilepsy [10] and Parkinson’s disease [11, 12, 13, 14]. These surrogate models have two broad applications: parameter inference and control. Parameter inference can help identify critical states and associated parameter values within a system, ideally also exposing underlying mechanistic processes. Control using surrogate models aims to exploit system knowledge to improve control performance. Methods such as model predictive control [15] and model-based reinforcement learning [16] are examples.
@@
Real NLONs are often high dimensional, partially observable, noisy, and involve complex intra-network interactions. As such, it can be extremely difficult to create accurate surrogate models. This challenging task may be approached in several ways. The classical approach is direct physics-based modeling, where mechanisms are pre-ordained and parameters fit to data. Machine learning (ML), or data-driven modeling, is an alternative approach which uses fully parameterized models and with parameters updated using learning algorithms. These two contrasting methods confer distinct benefits and drawbacks.
@@
Physics-based models are physically accurate within the bounds of the assumptions made in their construction. Similarly, within the domain of their training data, data-driven models perform well, however physical accuracy is not enforced. Failure when predicting out-of-domain is therefore common to the two approaches. Out-of-domain failure is guaranteed for physics-based models as there is no scope for adapting their structure or parameters. ML models however, can be updated online using new data to adapt to new situations, although this is in itself a challenging problem [17]. Physics-based models are inherently interpretable: each term generally has some understood physical meaning, or represents some physical laws or constraints. On the other hand, ML discards this prior knowledge in favor of complete parameterisation from observed data. The correspondence between the parameters and the physical system can be difficult to discern. When data-driven models have many parameters, they require large amounts of data. Physics-based models tend to have few parameters that need fitting to data, and therefore generally require less. Both methods can be computationally expensive as physics-based models rely on complex numerical schemes, and data-driven models often need to use extensive training algorithms for parameter updates.
@@
PIML is a recently-formulated approach for surrogate modeling and prediction applications [18]. It combines both physics-based and data-driven methods, in an attempt to make use of the best features of each approach. Its goal is to obtain physically constrained, robust, and interpretable models that capture both expert knowledge of dynamical processes, and the information that can be extracted from data obtained from sensing and recording devices. PIML models promise to be more data efficient as they do not require all of the dynamics to be learned from scratch, and they may also facilitate adaptivity through the use of machine learning. PIML-based control for NLONs may thus result in more robust, efficient and accurate controllers that are adaptable and generalisable.
@@
For PIML-based modeling of dynamical systems, it is natural to consider ML components with a time component or sequential nature. For instance, recurrent neural networks (RNNs) and their variants (LSTMs [19], GRUs [20]) are a common choice [21, 22, 23]. However, RC [24, 25] is a particularly promising alternative, due to its small size and simple training procedure. The simplicity of an RC may also offer a unique benefit for PIML parameter inference; while a large RNN can learn, or over learn, a complete model of the data without any input from the physics-based component, an RC has a limited capacity which may force the model to use the physics-based component, making it more interpretable. The small number of parameters used by an RC may also further enhance the low-data requirement conferred by the use of system knowledge. RCs are a restricted form of recurrent neural network (RNN), where learning only takes place in an external readout layer. Sequential data is passed into the reservoir via a fixed, random input weight matrix. An update rule then acts as a discrete non-linear map upon the internal state stored within the activations of the reservoir nodes. The result is a high-dimensional non-linear filter of the incoming data with a fading memory of past states. Scaling the weights of the input matrix and internal connectivity controls the extent to which past-state information is maintained in the hidden state and how much influence is exerted by the input data. To compute an output from the reservoir’s internal state, an output weight matrix is trained, often using regularized linear regression. The readout can be trained to perform classification of the reservoir state, and thus the input sequence, or regression to predict numerical features. The readout layer may also be trained to perform $n$-step-ahead prediction. When configured to predict the next step in a sequence, the reservoir may be run autoregressively with its output fed back in as the next input instance and thus used for time-series forecasting. This is the format we are considering here: using RCs for the prediction of dynamical system trajectories.
@@
In our particular implementation, the residual physics is an additional higher harmonic in the coupling term for the ground truth Kuramoto-like system: we give the hybrid RC the standard Kuramoto model without this addition. This bi-harmonic Kuramoto model [30] produces behaviors not accessible to the original Kuramoto model. For example, when clustering of the oscillators around a phase occurs in the standard Kuramoto model, there is only one cluster; with an extra harmonic term this need not be true. The residual physics task aims to replicate realistic control scenarios with incomplete knowledge of the system structure. This is an interesting challenge for the hybrid RC since exact knowledge of the ground truth non-linearities has previously been identified as being crucial for Kuramoto oscillator network attractor reconstruction when using the Next Generation Reservoir Computer [31, 32]. In a similar fashion to the parameter error task, we run this test across a range of hyper parameters, with four qualitatively different dynamical regimes. We then use this to inform a demonstrative grid-search optimization process simulating the development of a surrogate model for control applications.
@@ Methods
We present first an overview of the standard and hybrid RCs used in this study, followed by a shared procedure for ground truth data generation and testing in each task. We then describe the specific initialization and training methods, followed by details of the parameter error and residual physics tasks.
@@
A hybrid RC [29] (Fig. 3) is constructed out of a standard echo state network with the addition of an ODE model of the target system (the ‘expert ODE model’, Fig. 3, blue triangle). An auxiliary next-step prediction $\tilde{\mathbf{u}}_{t+1}$ is computed from the present state $\mathbf{u}_{t}$ by a numerical integrator. The integrator’s prediction is then concatenated with the present state, $[\tilde{\mathbf{u}}_{t+1};\mathbf{u}_{t}]$, before being passed into the reservoir as usual via the input weight matrix. The integrator’s prediction, $\tilde{\mathbf{u}}_{t+1}$, is also passed around the reservoir and concatenated with the reservoir internal states after the internal update, $[\tilde{\mathbf{u}}_{t+1};\mathbf{r}_{t+1}]$. Because of the concatenation of the usual data instances with the predictions of the ODE model, the size of the input and output weight matrices must change accordingly: $\mathbf{B}\in\mathbb{R}^{D_{r}\times 2D_{u}}$, and $\mathbf{C}\in\mathbb{R}^{D_{u}\times(D_{r}+D_{u})}$. If a non-linear transformation is applied, it is applied to the reservoir internal states, $\mathbf{r}_{t+1}$, only, not the augmented state vector as a whole.
@@
The spectral radius and input scaling were set to 0.4 and 0.15 respectively as baseline values as in previous work[29].
@@
where $\mathbf{u}^{*}(t)$ is the prediction and $\mathbf{u}(t)$ the ground truth. This metric was used in previous work[29] to evaluate the valid time metric. For each test, the mean NMSE across the entire trajectory was computed, capturing the overall agreement between the forecast and the ground truth.
@@
Throughout this task, we used $\epsilon=0.4$.
@@
To explore the potential of ‘optimally’ tuned standard and hybrid RCs, a grid search was conducted in the residual physics task. High and low parameter settings for the input scaling, spectral radius, and regularization strength were chosen from the results of the parameter sweeps. In particular, these were selected to approximately range the optimum region identified for each parameter, across both the standard and hybrid RC results. A set of all of the combinations of these parameters was produced, labeled from A to H, at each vertex of a cube in parameter space (Fig. 7). For each parameter combination, we followed the shared evaluation method as for both of the parameter sweeps. $40$ instantiations of standard and hybrid reservoirs were produced and then trained and tested on ground truth data from three of the four regimes in the residual physics task. These were the synchronous, heteroclinic cycles, and partial synchrony regimes. We excluded the asynchronous regime as both the standard and hybrid RC failed to predict the dynamics in this regime.
@@ Results
In this section we demonstrate the capability of the hybrid RC to predict NLONs in a range of different circumstances. We first present results from the parameter error task. Across three qualitatively different dynamical regimes, we compare the mean performance of multiple standard and hybrid RCs. As a control, we test the hybrid RC’s expert model, the base ODE model, which is also subject to parameter error. We show two parameter sweeps: the spectral radius and the input scaling. Second, we give the results of the residual physics task, where the hybrid RC is given a reduced form of the ground truth bi-harmonic Kuramoto model. Parameter sweeps are presented for four dynamical regimes. First, we test synchronous and asynchronous regimes. These were present in the standard Kuramoto model, however the bi-harmonic synchronous regime can support multiple clusters. The other two regimes are heteroclinic cycles, and partial synchrony; both are inaccessible to the standard Kuramoto model. We test the effects of spectral radius, input scaling, regularization strength and reservoir size. Short-term prediction quality is evaluated to investigate viability for control applications. Finally, the results of a grid search optimization process across three parameters are presented, illustrating the potential of a tuned hybrid RC.
@@
Our initial aim was to assess whether hybrid RCs are a viable architecture for surrogate modeling of NLONs by testing the quality of their long-term prediction across a range of parameters. We find that the performance of the hybrid RC is consistently better than the base ODE model or the standard RC. The hybrid RC also achieves low error even when both of its constituent parts, the base ODE model and the standard RC, do not. This matches what was previously reported for the Lorenz system [29]. Despite some variation due to parameter tuning, the hybrid RC often achieves near zero mean NMSE, particularly for the synchronous and multi-frequency regimes (Fig. 8 a, d, c, f).
@@
Figure 9: Residual physics task parameter sweeps evaluating the short-term forecast performance of the hybrid RC with missing non-linearities in its expert ODE model. Mean valid time, $t^{*}$, achieved by the hybrid RC (red), standard RC (blue), and the base ODE model (black) across the four dynamical regimes, and four different parameter sweeps. Column: Dynamical regime, synchronous (a, e, i, m), asynchronous (b, f, j, n), heteroclinic cycles (c, g, k, o), partial synchrony (d, h, l, p). Row: Parameter varied, spectral radius (a, b, c, d), input scaling (e, f, g, h), regularization (i, j, k, l), reservoir size, $D_{r}$, (m, n, o, p). Individual dots are individual reservoir/ODE instantiations, each representing the mean valid time across 20 forecasts. Solid lines are the mean of the mean valid time across the reservoir/ODE instantiations. Shaded regions are one standard deviation across reservoir/ODE instantiations. The hybrid RC generally outperforms the standard RC and the base ODE model even on regimes out of its expert model’s domain. The standard and hybrid RCs fail to learn the asynchronous regime. The spectral radius significantly affects performance, with the standard RC showing degradation above a spectral radius of $1.0$. The hybrid RC behaves differently, showing recovery of performance when the spectral radius is above $1.0$. Input scaling primarily affects the heteroclinic cycles and partial synchrony regimes, with optimum performance reached at minimal input scaling. Regularization strongly affects the performance of both the hybrid and standard RCs, with low regularization causing failure. Hybrid RCs have a broader range of viable regularization strength. Reservoir size has little effect as long as a minimum of 100 nodes are available.
@@
As in the parameter-error task we find that the hybrid RC generally performs better than either the base ODE model or the standard RC (Fig. 9), although this is more variable than in the parameter-error task. This is true even on the out of domain heteroclinic cycles and partial synchrony regimes not accessible by the expert model alone. Both the standard and hybrid RC failed to learn the asynchronous regime, achieving $0.0$ s valid time across nearly all parameter settings (Fig. 9 b, f, j, n). The base ODE model achieves higher valid times on the asynchronous regime, however it is clear that it is also not accurate as it predicts slow, synchronous trajectories; it fails to capture the qualitative character of the regime (Supplementary Fig. S11).
@@
The regularization parameter (Fig. 9: i, j, k, l) has a strong effect on the valid time of both the standard and hybrid RC. High regularization in the synchronous regime greatly reduces the performance of the standard RC. In the heteroclinic cycles and partial synchrony dynamical regimes, sufficiently low regularization appears to cause complete failure of both the standard and hybrid RC. In all cases, the optimal regularization strength range is much broader for the hybrid RC than the standard RC.
Beyond requiring at least $100$ nodes, the reservoir size (Fig 9 m, n, o, p) has little effect on the valid time on any of the dynamical regimes. The hybrid RC benefits slightly from a larger reservoir ($350$ nodes) when predicting the heteroclinic cycles regime but it still outperforms the standard RC across the full range of reservoir sizes explored.
@@
In the final part of this study we investigate how easy it is to optimize hybrid RCs. We do this with control applications in mind, as optimization is an important step in the development of controllers, and short-term prediction horizons are well suited to them. We evaluated the performance of hybrid and standard RCs using the valid time metric in a grid search of the hyper parameters that were shown to most strongly affect performance (Fig. 9). We ran tests on three of the dynamical regimes (synchronous, heteroclinic cycles and partial synchrony). On each regime, we evaluated the hybrid RC and standard RC with parameter settings defined by the corners of a cube in parameter space, labeled A to G (Fig. 7).

### v2

@@ Abstract
Surrogate modeling of non-linear oscillator networks remains challenging due to discrepancies between simplified analytical models and real-world complexity. To bridge this gap, we investigate hybrid reservoir computing, combining reservoir computing with “expert” analytical models. Simulating the absence of an exact model, we first test the surrogate models with parameter errors in their expert model. Second, in a residual physics task, we assess their performance when their expert model lacks key non-linear coupling terms present in an extended ground-truth model. We focus on short-term forecasting across diverse dynamical regimes, evaluating the use of these surrogates for control applications. We show that hybrid reservoir computers generally outperform standard reservoir computers and exhibit greater robustness to parameter tuning. This advantage is less pronounced in the residual physics task. Notably, unlike standard reservoir computers, the performance of the hybrid does not degrade when crossing an observed spectral radius threshold. Furthermore, there is good performance for dynamical regimes not accessible to the expert model, demonstrating the contribution of the reservoir.
@@ Introduction
Networks of oscillators appear widely across engineering, and in both the physical and biological sciences. When the networks are non-linear their dynamical behavior can be complex, displaying synchronization, chaos, and traveling waves [1, 2, 3]. Creating predictive dynamical models is important in areas such as surrogate modeling of non-linear oscillator networks (NLONs) for smart electrical grid optimization [4, 5, 6], biological computing [7], synthetic biology [8, 9] and the diagnosis and treatment of neurological disorders such as epilepsy [10] and Parkinson’s disease [11, 12, 13, 14]. These surrogate models have two broad applications: parameter inference and control. Parameter inference can help identify critical states and associated parameter values within a system, ideally also exposing underlying mechanistic processes. Control using surrogate models aims to exploit system knowledge to improve control performance. Methods such as model predictive control [15] and model-based reinforcement learning [16] are examples.
@@
Real NLONs are often high dimensional, partially observable, noisy, and involve complex intra-network interactions. As such, it can be extremely difficult to create accurate surrogate models. The classical approach to this task is direct physics-based modeling, where mechanisms are pre-ordained and parameters fit to data. Machine learning (ML), or data-driven modeling, is an alternative approach which uses fully parameterized models and with parameters updated using learning algorithms. These two contrasting methods confer distinct benefits and drawbacks.
@@
Physics-based models are physically accurate within the bounds of the assumptions made in their construction. Similarly, within the domain of their training data, data-driven models perform well. Failure when predicting out-of-domain is therefore common to the two approaches. ML models can be updated online using new data to adapt to new situations, although this is in itself a challenging problem [17]. Physics-based models are inherently interpretable: each term generally has some understood physical meaning, or represents some physical laws or constraints. On the other hand, ML discards this prior knowledge in favor of complete parameterisation from observed data. When data-driven models have many parameters, they require large amounts of data. Physics-based models tend to have few parameters that need fitting to data, and therefore generally require less. Both methods can be computationally expensive as physics-based models rely on complex numerical schemes, and data-driven models often need to use extensive training algorithms for parameter updates.
@@
PIML is a recently-formulated approach for surrogate modeling and prediction applications [18]. It combines both physics-based and data-driven methods, attempting to make use of the best features of each. Its goal is to obtain physically constrained, robust, and interpretable models that capture both expert knowledge of dynamical processes and the information that can be extracted from data obtained from sensing and recording devices. PIML models promise to be more data efficient as they do not require all of the dynamics to be learned from scratch, and they may also facilitate adaptivity through the use of machine learning. PIML-based control for NLONs may thus result in more robust, efficient and accurate controllers that are adaptable and generalisable.
@@
For PIML-based modeling of dynamical systems, it is natural to consider ML components with a time component or sequential nature. For instance, recurrent neural networks (RNNs) and their variants (LSTMs [19], GRUs [20]) are a common choice [21, 22, 23]. However, RC [24, 25] is a particularly promising alternative, due to its simple training procedure. The simplicity of an RC may also offer a unique benefit for PIML parameter inference; while a large RNN can learn, or over learn, a complete model of the data without any input from the physics-based component, an RC has a limited capacity which may force the model to use the physics-based component, making it more interpretable. The small number of parameters used by an RC may also further enhance the low-data requirement conferred by the use of system knowledge. RCs are a restricted form of recurrent neural network (RNN), where learning only takes place in an external readout layer. Sequential data is passed into the reservoir via a fixed, random input weight matrix. An update rule then acts as a discrete non-linear map upon the internal state stored within the activations of the reservoir nodes. The result is a high-dimensional non-linear filter of the incoming data with a fading memory of past states. Scaling the weights of the input matrix and internal connectivity controls the extent to which past-state information is maintained in the hidden state and how much influence is exerted by the input data. To compute an output from the reservoir’s internal state, an output weight matrix is trained, often using regularized linear regression. The readout layer may be trained to perform $n$-step-ahead prediction. When configured to predict the next step in a sequence, the reservoir may be run autoregressively with its output fed back in as the next input instance and thus used for time-series forecasting. This is the format we are considering here: using RCs for the prediction of dynamical system trajectories.
@@
In our implementation, the residual physics is an additional higher harmonic in the coupling term for the ground truth Kuramoto-like system: we give the hybrid RC the standard Kuramoto model without this addition. This bi-harmonic Kuramoto model [30] produces behaviors not accessible to the original Kuramoto model. For example, when clustering of the oscillators around a phase occurs in the standard Kuramoto model, there is only one cluster; with an extra harmonic term this need not be true. The residual physics task aims to replicate realistic control scenarios with incomplete knowledge of the system structure. This is an interesting challenge for the hybrid RC since exact knowledge of the ground truth non-linearities has previously been identified as being crucial for Kuramoto oscillator network attractor reconstruction when using the Next Generation Reservoir Computer [31, 32]. We run this test across a range of hyper parameters, with four qualitatively different dynamical regimes, using the results to inform a demonstrative grid-search optimization process simulating the development of a surrogate model for control applications. Overall, we are not seeking the best method for modeling NLONs [33, 34], instead, we look to understand if the benefit of the hybrid approach seen in  [29] carries over to application to NLONs and the more realistic residual physics task.
@@ Methods
We present first an overview of the standard and hybrid RCs used in this study, followed by a shared procedure for ground truth data generation and testing in each task. We then describe the specific initialization and training methods, followed by details of the parameter error and residual physics tasks. The details of the hybrid RC and parameter error task largely follow the approach in  [29]. They can therefore be skipped if the reader is familiar with this prior work. The only exception is our addition of the “Phase component transformation” to deal with the phase variables. Our aim is to test hybrid reservoir computing as an approach, rather than to tune a model to a particular application, so we have used as a baseline the parameter settings previously used in  [29]; this isolates the consequence of using the new ground truth system and the new residual physics task from the effect of parameter tuning.
@@
A hybrid RC [29] (Fig. 3) is constructed out of a standard echo state network with the addition of an ODE model of the target system (the ‘expert ODE model’, Fig. 3, blue triangle). An auxiliary next-step prediction $\tilde{\mathbf{u}}_{t+1}$ is computed from the present state $\mathbf{u}_{t}$ by a numerical integrator. The integrator’s prediction is then concatenated with the present state, $[\tilde{\mathbf{u}}_{t+1};\mathbf{u}_{t}]$, before being passed into the reservoir as usual via the input weight matrix. The integrator’s prediction, $\tilde{\mathbf{u}}_{t+1}$, is also passed around the reservoir and concatenated with the reservoir internal states after the internal update, $[\tilde{\mathbf{u}}_{t+1};\mathbf{r}_{t+1}]$. Because of the concatenation of the usual data instances with the predictions of the ODE model, the size of the input and output weight matrices must change accordingly: $\mathbf{B}\in\mathbb{R}^{D_{r}\times 2D_{u}}$, and $\mathbf{C}\in\mathbb{R}^{D_{u}\times(D_{r}+D_{u})}$. If a non-linear transformation is applied, it is applied to the reservoir internal states, $\mathbf{r}_{t+1}$, only, not the augmented state vector as a whole. The dual use of the integrator’s prediction $\tilde{\mathbf{u}}_{t+1}$ is present to maintain the same approach as  [29], despite it obscuring where the benefit of the hybrid approach originates.
@@
The spectral radius and input scaling were set to 0.4 and 0.15 respectively as baseline values as in previous work[29]. Although restricting the input to one per node may not result in optimum performance due to a lack of mixing at the present step, this was done to match the approach in  [29] to allow easy comparison of the results.
@@
where $\mathbf{u}^{*}(t)$ is the prediction and $\mathbf{u}(t)$ the ground truth. This metric was used in previous work[29] to evaluate the valid time metric. For each test, the mean NMSE across the entire trajectory was computed, capturing the overall, long-term, agreement between the forecast and the ground truth.
@@
Throughout this task, we used $\epsilon=0.4$. This metric only reports the short term accuracy, and therefore does  not provide information on the long term prediction quality of the models. For our application focus, where short time horizons are often all that is required, this provides a sufficient view of the performance.
@@
To explore the potential of ‘optimally’ tuned standard and hybrid RCs, a grid search was conducted in the residual physics task. We consider here only parameter tuning of the hybrid and standard RCs to maintain a consistent architecture and approach to  [29]. Prior works have explored optimization of the particular combination of model-based and data-driven components in the hybrid RC [38]. High and low parameter settings for the input scaling, spectral radius, and regularization strength were chosen from the results of the parameter sweeps. In particular, these were selected to approximately range the optimum region identified for each parameter, across both the standard and hybrid RC results. A set of all of the combinations of these parameters was produced, labeled from A to H, at each vertex of a cube in parameter space (Fig. 7). For each parameter combination, we followed the shared evaluation method as for both of the parameter sweeps. $40$ instantiations of standard and hybrid reservoirs were produced and then trained and tested on ground truth data from three of the four regimes in the residual physics task. These were the synchronous, heteroclinic cycles, and partial synchrony regimes.
@@ Results
In this section we demonstrate the capability of the hybrid RC to predict NLONs in various circumstances. We first present results from the parameter error task. Across three qualitatively different dynamical regimes, we compare the mean performance of multiple standard and hybrid RCs. As a control, we test the hybrid RC’s expert model, the base ODE model, which is also subject to parameter error. We show parameter sweeps of spectral radius and input scaling. Second, we give the results of the residual physics task, where the hybrid RC is given a reduced form of the ground truth bi-harmonic Kuramoto model. Parameter sweeps are presented for four dynamical regimes. First, we test synchronous and asynchronous regimes. These were present in the standard Kuramoto model, however the bi-harmonic synchronous regime can support multiple clusters. The other two regimes are heteroclinic cycles, and partial synchrony; both are inaccessible to the standard Kuramoto model. We test the effects of spectral radius, input scaling, regularization strength and reservoir size. Short-term prediction quality is evaluated to investigate viability for control applications. Finally, the results of a grid search optimization process across three parameters are presented, illustrating the potential of a tuned hybrid RC.
@@
Our initial aim was to assess whether hybrid RCs are a viable architecture for surrogate modeling of NLONs by testing the long-term prediction quality across a range of parameters. We find that the performance of the hybrid RC is consistently better than the base ODE model or the standard RC. The hybrid RC achieves low error even when both of its constituent parts, the base ODE model and the standard RC, do not. This matches what was previously reported for the Lorenz system [29]. Despite some variation due to parameter tuning, the hybrid RC often achieves near zero mean NMSE, particularly for the synchronous and multi-frequency regimes (Fig. 8 a, d, c, f).
@@
Figure 9: Residual physics task parameter sweeps evaluating the short-term forecast performance of the hybrid RC with missing non-linearities in its expert ODE model. Mean valid time, $t^{*}$, achieved by the hybrid RC (red), standard RC (blue), and the base ODE model (black) across the four dynamical regimes, and four different parameter sweeps. Column: Dynamical regime, synchronous (a, e, i, m), asynchronous (b, f, j, n), heteroclinic cycles (c, g, k, o), partial synchrony (d, h, l, p). Row: Parameter varied, spectral radius (a, b, c, d), input scaling (e, f, g, h), regularization (i, j, k, l), reservoir size, $D_{r}$, (m, n, o, p). Individual dots are individual reservoir/ODE instantiations, each representing the mean valid time across 20 forecasts. Solid lines are the mean of the mean valid time across the reservoir/ODE instantiations. Shaded regions are one standard deviation across reservoir/ODE instantiations. The hybrid RC generally outperforms the standard RC even on regimes out of its expert model’s domain. The spectral radius significantly affects performance, with the standard RC showing degradation above a spectral radius of $1.0$. The hybrid RC behaves differently, showing recovery of performance when the spectral radius is above $1.0$. Input scaling primarily affects the asynchronous, heteroclinic cycles and partial synchrony regimes, with optimum performance reached at minimal input scaling. Regularization strongly affects the performance of both the hybrid and standard RCs, with low regularization causing failure. Hybrid RCs have a broader range of viable regularization strength on the heteroclinic cycles and partial synchrony regimes. Reservoir size has little effect as long as a minimum of 100 nodes are available.
@@
As in the parameter-error task we find that the hybrid RC generally performs better than the standard RC (Fig. 9), although this is more variable than in the parameter-error task and the performance of the two is more similar. This is true even on the out of domain heteroclinic cycles and partial synchrony regimes not accessible by the expert model alone. In the asynchronous case, the standard RC outperforms the hybrid RC, except for cases with spectral radii above 1.0. The base ODE model achieves higher valid times than the standard or hybrid RC across the partial synchrony regime sweeps (Fig. 9 d, h, i, p), and occasionally for the asynchronous (Fig. 9 b, f, j, n) and heteroclinic cycles regimes (Fig. 9 c, g, k, o). It is clear however, that the long term prediction of the base ODE model is not accurate as it fails to capture the qualitative character of these regimes (Supplementary Figures  S11,  S12,  S13).
@@
The regularization parameter (Fig. 9: i, j, k, l) has a strong effect on the valid time of the standard and hybrid RC. High regularization in the synchronous regime greatly reduces the performance of the standard RC. In the asynchronous, heteroclinic cycles, and partial synchrony dynamical regimes, sufficiently low regularization appears to cause failure of the standard and hybrid RC. In all cases, the optimal regularization strength range is much broader for the hybrid RC than the standard RC.
Beyond requiring at least $100$ nodes, the reservoir size (Fig 9 m, n, o, p) has little effect on the valid time on any of the regimes. The hybrid RC benefits slightly from a larger reservoir ($350$ nodes) when predicting the heteroclinic cycles regime but it still outperforms the standard RC across the full range of reservoir sizes explored.
@@
In the final part of this study we investigate how easy it is to optimize hybrid RCs. We evaluated the performance of hybrid and standard RCs using the valid time metric in a grid search of the hyper parameters that were shown to most strongly affect performance (Fig. 9). We ran tests on three of the dynamical regimes (synchronous, heteroclinic cycles and partial synchrony), evaluating the hybrid RC and standard RC with parameter settings defined by the corners of a cube in parameter space, labeled A to G (Fig. 7).